OpenStudy (thecatman):

Charla and David are observing the leaf growth rates of their plants. Charla’s plant starts on day 1 with 43 leaves and gets 1 new leaf a day. David’s plant starts on day 1 with 27 leaves and gets 5 leaves a day. How many days will it take David’s plant to have as many leaves as Charla’s plant? 7 4 9 8

OpenStudy (thecatman):

b i think

OpenStudy (paki):

see.. Pc(Charla's plant)=42+n (where n is the number of days) Pd(David Plant) = 7n so we want to find when Pd=Pc Solving.. 7n=42+n 6n=42 n=7. So, on day 7, David's plant will have as many leaves as Charla's plant....